A roofer working on a 30° incline roof drops his hammer. the hammer slides 2.8 m with an acceleration of 1.3 m/s squared along the roof before falling towards the ground

if the hammer lands 3.2 m horizontally from where It left the roof how high is the lip of the roof?
Answer #1

hammer slides on roof from A to B and drops to point c. a veoro conto S=2.8 m - - - Vina - O = 30 - - Ground x=3.2m V- elocitApplying 1st equation of motion along the plane - V₂ = VA + at Applying 3 rd equation of motion along the incline- v²_v² = 2aV= 7.28 v=2.7 m/s vcoso = V C0830 = 2.7 x S3 2 = 2.33 mles veino = vain 30 2.7 x 1 = 1.35 mle from B to cin Applying 2nd equaApplying 2nd equation of motion in vertical direction from B to C- 2 h= Vot + 1 at? V = Vsino = 1.35 m/s a acceleration is ve